OpenStudy (anonymous):

convert 1.42 g/cm^2 to mg/mm^2 step by step?

OpenStudy (anonymous):

\[1.42 \frac{ g}{cm^2}(\frac{1000mg}{1 g})(\frac{1cm^2}{100 mm^2})=14.2\frac{mg}{mm^2}\]

OpenStudy (anonymous):

Dimensional analysis is your friend for this one.